You buy some beads and string from a student on campus to make a bracelet for yourself. You valued the materials at $10 but were charged only $8. Unknown to you, the seller would have sold them for $5. After you make the bracelet, you like it more than the original materials used to make it. You value the bracelet at $17. One day, a random stranger (who values your bracelet at $28) compliments the bracelet and offers you $20 for it, and you decide to sell it to the stranger. What is the total surplus generated from the original purchase of beads and string combined with the sale of the bracelet to the stranger? Total surplus = $
Answer #1

Let us consider the sale of beads first,

Consumer surplus=Willingness to pay-Price=10-8=$2

Producer surplus=Price-Willingness to sell=8-5=$3

Total surplus=CS+PS=2+3=$5

Let the consider the sale of bracelet

Consumer surplus=Willingness to pay-Price=28-20=$8

Producer surplus=Price-Willingness to sell=20-17=$3

Total surplus=CS+PS=8+3=$11

Total Surplus generated=5+11=$16